# Learning from core-collapse supernova remnants on the explosion mechanism

### v1

@@ Abstract
Based on the recently solidified notion that the jittering jets explosion mechanism (JJEM) is the primary explosion mechanism of core-collapse supernovae (CCSNe), I estimate some typical properties of the jittering jets. From the imprints of jittering jets in the outskirts of some CCSN remnants, I estimate the half-opening angles of jittering jets that shape CCSN remnants to be $\alpha_{\rm j}\simeq 1^{\circ}–10^{\circ}$. I also estimate that intermittent accretion disks around the newly born neutron star (NS) can launch jets after they live for only several times their orbital period around the NS. To operate, the JJEM requires the intermittent accretion disks that launch the jets to amplify the magnetic fields in a dynamo and the magnetic fields to reconnect to release their energy rapidly. I estimate the width of magnetic field reconnection zones to be $D_{\rm rec}\approx 0.005r\approx 0.1{~{}\rm km}$ near the surface of the NS. This width requires a numerical resolution several times smaller than the resolution of present CCSN simulations. I argue, therefore, that existing simulations of the CCSN explosion mechanism are still far from correctly simulating CCSN explosions.
@@ 1 Introduction
Two competing theoretical explosion mechanisms of massive stars utilize the gravitational energy the collapsing core releases as it forms a new neutron star (NS) to power core-collapse supernovae (CCSNe). In the delayed-neutrino explosion mechanism, neutrinos mediate the gravitational energy to explode the star (e.g., Bethe & Wilson 1985; Heger et al. 2003; Nordhaus et al. 2010, 2012; Janka 2012; Burrows 2013; Müller et al. 2019, 2024; Fujibayashi et al. 2021; Fryer et al. 2022; Boccioli et al. 2022, 2023; Boccioli & Roberti 2024; Nakamura et al. 2022, 2024; Olejak et al. 2022; Burrows et al. 2023; Andresen et al. 2024; Burrows et al. 2024b; Janka & Kresse 2024; Müller 2024; van Baal et al. 2024; Wang & Burrows 2024; Laplace et al. 2024). The magnetorotational explosion mechanism, where a pair of energetic opposite jets with a fixed axis explode the star, requires a rapidly rotating pre-collapse core (for recent studies see, e.g., Shibagaki et al. 2024; Zha et al. 2024a); therefore, it can power only a very small fraction of CCSNe. Despite being powered by jets, I take the magnetorotational explosion mechanism part of the neutrino-driven mechanism because it attributes most CCSNe, those with no pre-collapse rapidly rotating cores, to the delayed-neutrino explosion mechanism.
@@
The morphological signatures left by pairs of jets in many CCSN remnants (CCSNRs) are compatible with the JJEM, which considers such jets part of the explosion mechanism. Such signatures include opposite pairs of ears, nozzles, clumps, and filaments in CCSNRs, (e.g., Grichener & Soker 2017 who estimated the energies of the pairs of jets in several CCSNRs). The similarities of jet-shaped morphological features in CCSNRs with jet-shaped morphologies in planetary nebulae (e.g., Bear et al. 2017; Bear & Soker 2017, 2018; Soker 2022b, 2024a; Bear et al. 2024) and in cooling flow clusters (Soker, 2024b) solidify the claim for CCSNR shaping by energetic jets. The claim for shaping by energetic jets that are part of the pairs of jets that explode progenitors of the CCSNRs is particularly strong on CCSNRs that possess point-symmetric morphology (Soker & Shishkin, 2024). A point-symmetric morphology has two or more pairs of opposite structural features (ears, clumps, filaments, nozzles) that do not share the same symmetry axis. The present list of point-symmetric CCSNRs and the studies that attributed their morphologies to jittering jets are: SNR 0540-69.3 (Soker, 2022a), the Vela CCSNR (Soker 2023b; Soker & Shishkin 2024), CTB 1 (Bear & Soker, 2023), N63A (Soker, 2024c), SN 1987A (Soker, 2024d, e), G321.3–3.9 (Soker, 2024b; Shishkin & Soker, 2024), G107.7-5.1 (Soker, 2024b), Cassiopeia A (Bear & Soker, 2024), the Cygnus Loop (Shishkin et al., 2024), and Puppis A (Bear et al., 2024).
Based on this list of point-symmetric CCSNRs and other CCSNRs with signatures of one pair of jets, I consider the JJEM to be CCSNe’s primary or even sole explosion mechanism. This implies that the newly born NS manages to launch jets within several seconds from the formation of the proto-NS and the bounce of the shock that the proto-NS forms as the collapse stops due to nuclear forces. I use this implication to deduce some properties of the jet-launching process from young NSs (Section 2). I then show (Section 3) that to simulate the relevant processes for the JJEM, three-dimensional (3D) magneto-hydrodynamical (MHD) numerical codes must have very high spatial resolution beyond the reach of most present computers available to the community. I summarize this short study and conclude that the community is far from simulating the correct CCSN explosion process.
@@ 2.1 Previously derived JJEM parameters
According to the JJEM, the basic explosion process starts when the newly born NS (and later the black hole if the NS collapses into a black hole) accretes mass with stochastic angular momentum from the collapsing core via intermittent accretion disks. The stochastic angular momentum results from instabilities above the NS, including the spiral standing accretion shock instability (e.g., Buellet et al. 2023 for a recent study of this instability), that amplify angular momentum seed fluctuations in the collapsing core material. The seed fluctuations come from the convective motion in the pre-collapse core (e.g., Papish & Soker 2014a; Gilkis & Soker 2014, 2016; Shishkin & Soker 2023; Wang et al. 2024), and, in some cases, possibly from the envelope (e.g., Quataert et al. 2019; Antoni & Quataert 2022, 2023). In cases of a rapid pre-collapse core rotation, the stochastic angular momentum variations are around this rotation axis (e.g., Soker 2023a). Namely, it is not entirely stochastic at a whole solid angle.
Earlier studies have derived some of the basic properties of the JJEM, which I summarize in Table 1. The first 11 properties are in a table from an earlier paper (Soker, 2024e). In this study, I added the last three rows (12-14). Table 1 presents the range of typical values of the JJEM and the justifications for these values. The values the table lists are for iron CCSNe, not electron-capture CCSNe. In electron-capture CCSNe, the primary explosion phase (jet-launching period) might be much longer, up to several minutes to even a few hours rather than seconds (Wang et al., 2024).
@@
# The values in the table are for CCSNe of collapsing iron-rich core. Jets of electron capture supernovae are less energetic, and the explosion might last for minutes to a few hours (Wang et al., 2024).
@@
References: [B+25]: Bear et al. (2024); [BGS17]: Bear et al. (2017); [GiSo15]: Gilkis & Soker (2015);[GrSo17]: Grichener & Soker (2017); [Li04]: Livio (2004); [Ni18]: Nisini et al. (2018); [PaSo11]: Papish & Soker (2011); [PaSo14a]: Papish & Soker (2014a); [PaSo14b]: Papish & Soker (2014b); [ShSo14]: Shishkin & Soker (2021); [So24e]: Soker (2024e); [So24f]: Soker (2024f).
@@ 2.2 Half-opening angle of the jets
In an earlier paper (Soker, 2022c), I suggested the collimation of the jets by the pressure of the ambient gas close to the NS, below the gain region, because the pressure in that region is much larger than in the gain region (e.g., Janka 2001). Further out, the material that the jet shocks, the so-called cocoon, maintains the collimation of the jet or might even compress it somewhat (Soker, 2022c). Here, I quantify the collimation and find that these collimation processes should be efficient in bringing the opening angle of the jets to be small.
Clumps are also observed in point-symmetric CCSNRs. Clumps can be the tip of jets. Another possibility is that jet-inflated bubbles in the core compress clumps, as Papish & Soker (2014a) have shown, and the clumps penetrate the ejecta and imprint the morphology. Similarities with cooling flow clusters support this process of forming clumps in some, but not all, cases (Soker, 2024b).
@@ 2.3 Minimum time for jet launching
Bear et al. (2024) proposed a new mechanism to impart a natal kick velocity to the NS, the kick by early asymmetrical pair (kick-BEAP) mechanism. In the kick-BEAP mechanism, the NS launches a pair (or two) of two opposite jets where one is much more powerful. The momentum asymmetry between the two jets is significant enough to impart a large velocity to the NS (Bear et al., 2024). The kick-BEAP operates within $t_{\rm b}\lesssim 0.2{~{}\rm s}$ after shock bounce when the accretion of core material onto the NS is very high. At that early phase, the radius of the hot NS is much larger than its final radius, $R(<0.2{~{}\rm s})\simeq 40{~{}\rm km}$ (e.g., Raynaud et al. 2020) instead of $R_{\rm Ns}\simeq 12{~{}\rm km}$; it is a proto-NS.
The jets of the kick-BEAP mechanism are launched from a radius of $\simeq 40{~{}\rm km}$. The orbital period at this radius, for a mass of still $M\simeq 1.2M_{\odot}$, is $\tau_{\rm Kep}\simeq 0.004{~{}\rm s}$. The high-mass accretion rate lasts for $\simeq 0.2{~{}\rm s}$ (e.g., Müller et al. 2017; Burrows et al. 2024a), and a jet-launching episode for a shorter time; Bear et al. (2024) scale with a jet launching period of $\Delta t_{1}=0.05{~{}\rm s}$. But the accretion disk starts to launch jets before it ceases to exist. Namely when its age is $<0.05{~{}\rm s}\simeq 10\tau_{\rm Kep}$.
@@ 2.4 The rotational energy of the accreted material
The orbital velocity near the surface of the NS during the explosion process when the NS is not yet settled to its final radius is $v_{\rm Kep}\simeq 10^{5}{~{}\rm km}{~{}\rm s}^{-1}$. This is about the typical jets’ velocity in the JJEM, $v_{\rm j}\simeq v_{\rm Kep}$ (Table 1). The two jets in each jet-launching episode carry a fraction $\eta_{\rm 2j}\simeq 0.1$ of the mass in the accretion disk so that a fraction of $1-\eta_{\rm 2j}$ is accreted. The jets carry angular momentum, and the accreted mass settles onto the NS with lower rotational velocity than in the accretion disk. As with young stellar objects that rotate at tens of percent of the break-up velocity (e.g., Huang et al. 2010), I take here the rotational velocity of the accreted mass to be $v_{\rm rot}=\delta_{\rm K}v_{\rm Kep}$, with $\delta_{\rm rot}\simeq 0.3-0.5$. Since $v_{\rm j}\simeq v_{\rm Kep}$, I take $v_{\rm rot}=\delta_{\rm j}v_{\rm j}$, with $\delta_{\rm j}\simeq 0.3-0.5$. The ratio of the rotational energy of the accreted mass in an episode to the energy of the jets of that episode is then
@@
Future magnetohydrodynamic simulations of the CCSN explosion process should include this friction between accreted material from consecutive jet-launching episodes. This demands a high, but not very high, numerical resolution. The zone between consecutive accretion episodes behaves like a boundary layer where the velocity has a large gradient. A boundary layer in fixed-axis accretion disks can boost mass ejection if shocks occur within it (e.g., Soker & Regev 2003). The width of the boundary layer is of the order of the disk’s half-thickness. Allowing for a disk half-thickness of $\simeq 0.1r$ and demanding resolving the boundary layer by at least four cells requires the numerical resolution near the NS to have a cell size of $\Delta\lesssim 0.025r$. This is about the resolution of present simulations of CCSNe, as I discuss in Section 3, where I also show that magnetic processes require even finer resolution.

### v2

@@ Abstract
I estimate some typical properties of the jittering jets explosion mechanism (JJEM) to distinguish it from competing supernova explosion models. From the imprints of jittering jets in the outskirts of some CCSN remnants, I estimate the half-opening angles of jittering jets that shape CCSN remnants to be $\alpha_{\rm j}\simeq 1^{\circ}-10^{\circ}$. I also estimate that intermittent accretion disks around the newly born neutron star (NS) can launch jets after they live for only several times their orbital period around the NS. To operate, the JJEM requires intermittent accretion disks that launch jets to amplify the magnetic fields in a dynamo, and the magnetic fields to reconnect and release their energy rapidly. I estimate the width of magnetic field reconnection zones to be $D_{\rm rec}\approx 0.005r\approx 0.1{~{}\rm km}$ near the surface of the NS. This width requires a numerical resolution several times smaller than the resolution of present CCSN simulations. I argue, therefore, that existing simulations of the CCSN explosion mechanism are still far from correctly simulating CCSN explosions.
@@ 1 Introduction
Two competing theoretical explosion mechanisms of massive stars utilize the gravitational energy released by the collapsing core as it forms a new neutron star (NS) to power core-collapse supernovae (CCSNe). In the delayed-neutrino explosion mechanism, neutrinos mediate the gravitational energy to explode the star (e.g., Bethe & Wilson 1985; Heger et al. 2003; Nordhaus et al. 2010, 2012; Janka 2012; Burrows 2013; Müller et al. 2019, 2024; Fujibayashi et al. 2021; Fryer et al. 2022; Boccioli et al. 2022, 2023; Boccioli & Roberti 2024; Nakamura et al. 2022, 2025; Olejak et al. 2022; Burrows et al. 2023; Andresen et al. 2024; Burrows et al. 2024b; Janka & Kresse 2024; Müller 2024; van Baal et al. 2024; Wang & Burrows 2024; Laplace et al. 2024; Huang et al. 2024). The magnetorotational explosion mechanism, where a pair of energetic opposite jets with a fixed axis explode the star, requires a rapidly rotating pre-collapse core (for recent studies see, e.g., Shibagaki et al. 2024; Zha et al. 2024a); therefore, it can power only a very small fraction of CCSNe. Despite being powered by jets, I take the magnetorotational explosion mechanism part of the neutrino-driven mechanism because it attributes most CCSNe, those with no pre-collapse rapidly rotating cores, to the delayed-neutrino explosion mechanism.
@@
One of the largest achievements of simulations of the delayed-neutrino explosion mechanism is determining that heating by neutrinos is significant. Many such simulations manage to explode stellar models solely through neutrino heating. However, in many cases, no explosion occurs in the simulations. One of the disadvantages of the JJEM is that no simulations exist. As I demonstrate in this study, simulating the JJEM requires much more sophisticated numerical codes, which do not currently exist.
The morphological signatures left by pairs of jets in many CCSN remnants (CCSNRs) are compatible with the JJEM, which considers such jets part of the explosion mechanism. Such signatures include opposite pairs of ears, nozzles, clumps, and filaments in CCSNRs (e.g., Grichener & Soker 2017 who estimated the energies of the pairs of jets in several CCSNRs). The similarities of jet-shaped morphological features in CCSNRs with jet-shaped morphologies in planetary nebulae (e.g., Bear et al. 2017; Bear & Soker 2017, 2018; Soker 2022b, 2024a; Bear et al. 2025) and in cooling flow clusters (Soker, 2024b) solidify the claim for CCSNR shaping by energetic jets. The claim for shaping by energetic jets that are part of the pairs of jets that explode progenitors of the CCSNRs is particularly strong on CCSNRs that possess point-symmetric morphologies (Soker & Shishkin, 2025). A point-symmetric morphology has two or more pairs of opposite structural features (ears, clumps, filaments, nozzles) that do not share the same symmetry axis. The present list of point-symmetric CCSNRs and the studies that attributed their morphologies to jittering jets are: SNR 0540-69.3 (Soker, 2022a), CTB 1 (Bear & Soker, 2023), the Vela CCSNR (Soker 2023b; Soker & Shishkin 2025), N63A (Soker, 2024c), the Cygnus Loop (Shishkin et al., 2024), G321.3–3.9 (Soker, 2024b; Shishkin & Soker, 2025), SN 1987A (Soker, 2024d, e), G107.7-5.1 (Soker, 2024b), W44 (Soker, 2024f), Cassiopeia A (Bear & Soker, 2025), Puppis A (Bear et al., 2025), SNR G0.9+0.1 (Soker, 2025a), the Crab Nebula (Shishkin & Soker, 2024), N132D (Soker 2025b), and S147 Shishkin et al. 2025.
Based on this list of point-symmetric CCSNRs and other CCSNRs with signatures of one pair of jets, I consider the JJEM to be CCSNe’s primary or even sole explosion mechanism. This implies that the newly born NS manages to launch jets within several seconds from the formation of the proto-NS and the bounce of the shock that the proto-NS forms as the collapse stops due to nuclear forces. I use this implication to deduce some properties of the jet-launching process from young NSs (Section 2). I then demonstrate (Section 4) that to simulate the relevant processes for the JJEM, three-dimensional (3D) magnetohydrodynamic (MHD) numerical codes require a very high spatial resolution, which is beyond the reach of most present computers available to the community. I summarize this short study and conclude that the community is far from simulating the correct CCSN explosion process.
@@ 2.1 Previously derived JJEM parameters
According to the JJEM, the basic explosion process starts when the newly born NS (and later the black hole if the NS collapses into a black hole) accretes mass with stochastic angular momentum from the collapsing core via intermittent accretion disks. The stochastic angular momentum results from instabilities above the NS, including the spiral standing accretion shock instability (e.g., Buellet et al. 2023 for a recent study of this instability), that amplify angular momentum seed fluctuations in the collapsing core material. The seed fluctuations come from the convective motion in the pre-collapse core (e.g., Papish & Soker 2014a; Gilkis & Soker 2014, 2016; Shishkin & Soker 2023; Wang et al. 2024), and, in some cases, possibly from the envelope (e.g., Quataert et al. 2019; Antoni & Quataert 2022, 2023). In cases of a rapid pre-collapse core rotation, the stochastic angular momentum variations are around this rotation axis (e.g., Soker 2023a). Namely, it is not entirely stochastic over the entire solid angle.
I emphasize the ingredient of the JJEM that makes it unique and distinguishes it from the magnetorotational explosion mechanism. The magnetorotational explosion mechanism (e.g., Shibata et al. 2025) requires a rapidly pre-collapse rotating core; hence, it is rare. In the JJEM, the angular momentum of the intermittent accretion disks starts with angular momentum fluctuation in the convective zones of the core that are amplified by instabilities behind the stalled shock. The JJEM requires no pre-collapse core rotation. As I discuss in Section 4, current numerical simulations are unable to handle the necessary physics.
Earlier studies have derived some of the basic properties of the JJEM, which I summarize in Table 1. The first 11 properties are in a table from an earlier paper (Soker, 2024e). In this study, I added the last three rows (12-14). Table 1 presents the range of typical values of the JJEM and the justifications for these values. The values listed in the table are for iron CCSNe, not electron-capture CCSNe. In electron-capture CCSNe, the primary explosion phase (jet-launching period) might be significantly longer, lasting several minutes to even a few hours rather than seconds (Wang et al., 2024).
@@
# The values in the table are for CCSNe of collapsing iron-rich cores. Jets of electron-capture supernovae are less energetic, and the explosion might last for minutes to a few hours (Wang et al., 2024).
@@
References: [B+25]: Bear et al. (2025); [BGS17]: Bear et al. (2017); [GiSo15]: Gilkis & Soker (2015);[GrSo17]: Grichener & Soker (2017); [Li04]: Livio (2004); [Ni18]: Nisini et al. (2018); [PaSo11]: Papish & Soker (2011); [PaSo14a]: Papish & Soker (2014a); [PaSo14b]: Papish & Soker (2014b); [ShSo14]: Shishkin & Soker (2021); [So24e]: Soker (2024e); [So24f]: Soker (2024g).
@@ 2.2 Half-opening angle of the jets
In an earlier paper (Soker, 2022c), I suggested the collimation of the jets by the pressure of the ambient gas close to the NS, below the gain region, because the pressure in that region is much larger than in the gain region (e.g., Janka 2001). Further out, the material that the jet shocks, known as the cocoon, maintains the collimation of the jet or may even compress it somewhat (Soker, 2022c). Here, I quantify the collimation and find that these collimation processes should be efficient in reducing the opening angle of the jets to a small value.
Clumps are also observed in point-symmetric CCSNRs. Clumps can be the tip of jets, as in the 3D JJEM simulations by Braudo et al. (2025). Another possibility is that jet-inflated bubbles in the core compress clumps, as Papish & Soker (2014a) and Braudo et al. (2025) have shown, and the clumps penetrate the ejecta and imprint the morphology. The number of pairs of clumps can be larger than the number of pairs of jets (Braudo et al., 2025). Similarities with cooling flow clusters support this process of forming clumps in some, but not all, cases (Soker, 2024b).
@@ 2.3 Minimum time for jet launching
Bear et al. (2025) proposed a new mechanism to impart a natal kick velocity to the NS, the kick by early asymmetrical pair (kick-BEAP) mechanism. In the kick-BEAP mechanism, the NS launches a pair (or two) of opposite jets, where one jet is significantly more powerful. The momentum asymmetry between the two jets is significant enough to impart a large natal kick velocity to the NS (Bear et al., 2025). The kick-BEAP operates within $t_{\rm b}\lesssim 0.2{~{}\rm s}$ after shock bounce when the accretion of core material onto the NS is very high. At that early phase, the radius of the hot NS is much larger than its final radius, $R(<0.2{~{}\rm s})\simeq 40{~{}\rm km}$ (e.g., Raynaud et al. 2020) instead of $R_{\rm Ns}\simeq 12{~{}\rm km}$; it is a proto-NS.
The disk launches the jets of the kick-BEAP mechanism from a radius of $\simeq 40{~{}\rm km}$. The orbital period at this radius, for a mass of still $M\simeq 1.2M_{\odot}$, is $\tau_{\rm Kep}\simeq 0.004{~{}\rm s}$. The high-mass accretion rate lasts for $\simeq 0.2{~{}\rm s}$ (e.g., Müller et al. 2017; Burrows et al. 2024a), and a jet-launching episode for a shorter time; Bear et al. (2025) scale with a jet launching period of $\Delta t_{1}=0.05{~{}\rm s}$. But the accretion disk starts to launch jets before it ceases to exist. Namely when its age is $<0.05{~{}\rm s}\simeq 10\tau_{\rm Kep}$.
@@ 2.4 The rotational energy of the accreted material
The orbital velocity near the surface of the NS during the explosion process when the NS is not yet settled to its final radius is $v_{\rm Kep}\simeq 10^{5}{~{}\rm km}{~{}\rm s}^{-1}$. This is about the typical jets’ velocity in the JJEM, $v_{\rm j}\simeq v_{\rm Kep}$ (Table 1). The two jets in each jet-launching episode carry a fraction $\eta_{\rm 2j}\simeq 0.1$ of the mass in the accretion disk so that a fraction of $1-\eta_{\rm 2j}$ is accreted. The jets carry angular momentum, and the accreted mass settles onto the NS with lower rotational velocity than in the accretion disk. As with young stellar objects that rotate at tens of percent of the break-up velocity (e.g., Huang et al. 2010), I take here the rotational velocity of the accreted mass to be $v_{\rm rot}=\delta_{\rm K}v_{\rm Kep}$, with $\delta_{\rm K}\simeq 0.3-0.5$. Since $v_{\rm j}\simeq v_{\rm Kep}$, I take $v_{\rm rot}=\delta_{\rm j}v_{\rm j}$, with $\delta_{\rm j}\simeq 0.3-0.5$. The ratio of the rotational energy of the accreted mass in an episode to the energy of the jets of that episode is then
@@
Future magnetohydrodynamic simulations of the CCSN explosion process should include this friction between accreted material from consecutive jet-launching episodes. This demands a high, but not very high, numerical resolution. The zone between consecutive accretion episodes behaves like a boundary layer where the velocity has a large gradient. A boundary layer in fixed-axis accretion disks can boost mass ejection if shocks occur within it (e.g., Soker & Regev 2003). The width of the boundary layer is of the order of the disk’s half-thickness. Allowing for a disk half-thickness of $\simeq 0.1r$ and demanding resolving the boundary layer by at least four cells requires the numerical resolution near the NS to have a cell size of $\Delta\lesssim 0.025r$. This is about the resolution of current simulations of CCSNe, as I discuss in Section 4, where I also demonstrate that magnetic processes necessitate even finer resolution.
3 Observational properties
In Table 2, I compare the predicted observations of the JJEM with those of the neutrino-driven mechanism. A sign of (=)(=)( = ) indicates that the observational property in this row cannot support one explosion mechanism over the other. A sign of ()(-)( - ) means that the prediction of the explosion mechanism contradicts the considered observational property, while ()(--)( - - ) indicates severe contradiction. A sign of (+)(+)( + ) means that the explosion mechanism does better than the other mechanism in accounting for the observational property of the row. At the same time, (++)(++)( + + ) indicates that it does much better, to the degree of ruling out the other mechanism. Some rows in Table 2 contain all the necessary information; for others, I provide further elaboration below. In some places in the table, I list specific references; for others, the list of papers on the two explosion mechanisms in Section 1 contains the relevant information.
Table 2: Observational properties according to the two alternative explosion mechanisms
Observation Neutrino-driven JJEM
Lightcurve and spectrum (=)(=)( = ) For the same progenitor and the explosion energy, the JJEM predicts the same light curves and spectrum. The energy deposition in the inner zones, neutrino heating or jets, does not change much the emission from the outer regions.
NS mass (=)(=)( = ) Might be somewhat higher than the observed range [Sok24R] (=)(=)( = ) Determined by the location of strong convection in the pre-collapse core, gives the observed general range [ShSo22]
Black hole formation (=)(=)( = ) When neutrino heating does not revive the stalled shock, or does so only after large mass collapses to the center (=)(=)( = ) A rapidly rotating pre-collapse core leads to fixed-axis jets in the explosion, inefficiently ejecting stellar material from the equatorial plane
No failed supernovae ()(-)( - ) Most studies predict that a large fraction of massive stars do not explode, i.e., failed CCSNe (+)(+)( + ) No failed CCSNe as some collapsing gas has the angular momentum to form accretion disks that launch jets
Polarization (=)(=)( = ) Polarization results from instabilities, which can be significant and change direction with time. (=)(=)( = ) Polarization results from instabilities and jets that can change direction with time.
Nucleosynthesis yields (=)(=)( = ) The total mass of nucleosynthesis in the JJEM is as in the neutrino-driven mechanism for the same progenitor and explosion energy.
Nucleosynthesis morphologies (=)(=)( = ) Random clumps of heavy elements due to instabilities. (+)(+)( + ) Random clumps + metals might be in jet-shaped morphologies, e.g., S-rich jet in Cassiopeia A; S-shaped metals in the Vela SNR [SoSh25].
Neutrino emission (=)(=)( = ) Most neutrinos and all antineutrinos are from the cooling proto-NS. This is the same in the two mechanisms
[…]
NS kick and metals distribution (=)(=)( = ) Mainly asymmetrical mass ejection with opposite NS kick [HwLa12], e.g., the tugboat process [Sc04; No10] (=)(=)( = ) The tugboat process + the kick by early asymmetrical pairs (kick-BEAP) of jets [Be25], [Sh25]
Alignment of NS kick and spin ()(-)( - ) No explanation (+)(+)( + ) The kick-BEAP has the jets along the angular momentum of the accreted mass, and in some cases leads to spin-kick alignment [Be25]
Explosion energy of $E_{\rm exp}>2\times 10^{51}{~{}\rm erg}$ ()(-)( - ) Neutrino heating does not explain. Many consider a magnetar, but in many magnetar models the explosion itself is $E_{\rm exp}\gtrsim 3\times 10^{51}{~{}\rm erg}$, implying explosion by jets [SoGi17] [Ku25] (+)(+)( + ) Can explain all observed energies.
Point-symmetric CCSNRs ()(--)( - - ) Predicted morphological features, like ears, bays, and clumps, are generally randomly distributed and cannot explain point symmetry (++)(++)( + + ) Some of the pairs of jets that explode the star might leave point-symmetric morphological features
Notes: Comparing observational predictions of the neutrino-driven mechanism and the JJEM. Compatibility marks: (=)(=)( = ): observational property cannot be used to support one explosion mechanism over the other; ()(-)( - ): contradiction to observation; ()(--)( - - ): severe contradiction to observations; (+)(+)( + ): the mechanism does better than the other mechanism; (++)(++)( + + ): strongly support the mechanism as the primary explosion mechanism.
References: [Be25]: Bear et al. (2025); [HwLa12]: Hwang & Laming (2012)); [Ku25]: Kumar (2025); [No10]: Nordhaus et al. (2010); [Ma23]: Mezzacappa et al. 2023; [Sc04]: Scheck et al. (2004); [Sh25]: Shishkin et al. (2025); [ShSo22]: Shishkin & Soker (2022); [So23]: Soker (2023c); [Sok24R]: Soker (2024h); [SoGi17]: Soker & Gilkis (2017); [SoSh25]: Soker & Shishkin (2025);
Some of the outcomes of the processes occurring near the center of the star are similar in the two explosion mechanisms. One such process is the cooling of the proto-NS, which yields very similar, or even identical, neutrino emissions in the two explosion mechanisms I compare in Table 2. In the JJEM, the jets deposit energy in the center, rather than reviving the stalled shock. But once there are shocks in the core, the energy source does not matter for nucleosynthesis and energy deposition in the outer stellar zones. Therefore, the nucleosynthesis in the JJEM and the light curve and spectrum are similar to those calculated and simulated in the neutrino-driven mechanism. The large advantages of the neutrino-driven mechanism are that many simulations of these processes exist (references in Section 1). At the same time, there are no simulations yet from the collapse to the explosion of the JJEM (Section 4). One difference is in the distribution of the ejected metals from the core. The JJEM predicts that in some cases, the distribution will be shaped by jets, such as the sulfur-rich jet in Cassiopeia A and the S-shaped metal distribution in the Vela SNR that Soker & Shishkin (2025) attributed to the JJEM.
Most studies of the neutrino-driven mechanism predict that some massive stars fail to explode. This contradicts recent observations and their analysis, which imply that failed CCSNe are either rare or nonexistent (e.g., Byrne & Fraser 2022; Strotjohann et al. 2024; Beasor et al. 2025; Healy et al. 2024). Although Boccioli et al. (2025) suggested that the neutrino-driven mechanism can overcome this difficulty, due to the common view among its supporters, I indicate it as a problem for the neutrino-driven mechanism. According to the JJEM, an explosion, possibly an energetic one, occurs even when the remnant is a black hole (e.g., Gilkis et al. 2016).
Some pulsars have the NS spin and kick velocity aligned, or almost aligned (e.g., Johnston et al. 2005; Noutsos et al. 2012; Biryukov & Beskin 2024). The delayed neutrino mechanism has no explanation for that. A popular kick process in the neutrino-driven mechanism is the tugboat mechanism, where an ejected massive clump pulls the NS (e.g., Scheck et al. 2004, 2006; Nordhaus et al. 2010, 2012; Wongwathanarat et al. 2010, 2013; Janka 2017). In the JJEM, the NS acquires its natal kick velocity through the tugboat process and/or the newly proposed kick by early asymmetrical pairs (kick-BEAP) of jets (Bear et al., 2025). In the kick-BEAP process, one jet in a pair is much more powerful than the opposite jets; hence, momentum conservation implies an NS kick in the opposite direction of the more powerful jet. The jets expand along the angular momentum of the accretion disk that launches the jets, which is the kick direction in the kick-BEAP. If this accretion episode dominates the final angular momentum of the NS, then there is a spin-kick alignment.
The JJEM does better than the neutrino-driven mechanism in explaining no failed CCSNe, spin-kick alignment in some pulsars, and explosion energies of $E_{\rm exp}\gtrsim 3\times 10^{51}{~{}\rm erg}$. However, although these observations challenge the neutrino-driven mechanism, they do not rule it out. The most severe challenge, which seems to rule out the neutrino-driven mechanism as the primary explosion mechanism, is the identification of 15 point-symmetric CCSNRs (Section 1). The point-symmetric morphology in many, but not all, CCSNRs is a strong prediction of the JJEM, for which the neutrino-driven process has no explanation (see the discussion of this process by Soker & Shishkin 2025).